Astronomy – February 19th

Astronomy Report

Name: Benoit Floquet Crew: 189
date: 2018/02/19

Sky Conditions: The sky was clear during the window of observation. Before and after that it was cloudy

Wind Conditions: Quite a lot of wind but not that much during the window of observation.

Observation Start Time: 03:00 pm

Observation End Time: 03:30 pm

Summary: I went to the Musk Observatory to get a first experience with the telescope. I followed the instructions with success until I had to take a look at the Sun. It was not well aligned with the Sun Tracker but I couldn’t succeed to use the Hand control to make it right. Then I changed the set up ( time ) to get the Sun aligned and it was already to cloudy to take a good look…

Objects Viewed: none

Problems Encountered: I didn’t find the instructions to use the Hand control in order to move the telescope.

Operations Report – February 19th

Hi CapCom,

I’m Gabriel, and here is my Operation report.

Crew 189 Operations Report 19Feb2018

SOL: 1

Name of person filing report: Gabriel PAYEN

Non-nominal systems: Roof heater, pipe leak in low deck, ethernet hub, fan of backpack 4

Notes on non-nominal systems:

-Ethernet Hub :

This afternoon an experiment had trouble with an ethernet hub. We’ll try to fix it tonight or tomorrow.

-Fan :

During the engineering check all backpack seemed to be nominal but a fan on backpack 4 turned out of service during the EVA. Tomorrow the crew will be very busy, but we are planning an important backpack fixing operation during the upcoming days.

-Pipe leak :

The flow seems to be reduced, the bucket was empty this morning and at 50% tonight.

-Roof heater :

This morning it was very difficult to control the roof heater. Sometimes the flow became tepid or even cold without any reason because the furnace stops running. I’ve turned on the wall heater which runs all day.

Generator (hours run): All the day (turned on since 6:00 PM 18Feb) because of a cloudy day

Solar— SOC % (Before generator is run at night) 83% at 9 AM

Diesel – 70%

Propane – ?

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons.

Water (trailer) – 0 gallons.

Water (static) – 400 gallons (checked during the EVA).

Trailer to Static Pump used – No but Atila gave a 100 gallons load

Water (loft) – Static to Loft Pump used – YES

Water Meter: 0.24m

Toilet tank emptied: NO

ATV’s Used: None

Deimos rover used: YES

Hours: 109.1

Beginning charge: 100%

Ending charge: 93%

Currently charging: YES

Spirit rover used: NO

Hours: 20.3

Beginning charge: 100%

Ending charge: 100%

Currently charging:

Opportunity rover used: YES

Hours: 15.7

Beginning charge: 100%

Ending charge: 85%

Currently charging: YES

Curiosity rover used: NO

Hours: 12.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: NO

General notes and comments: Very busy today with my own experiments. As soon as possible (from sol 3 I hope) I’ll begin to deal with non-nominal systems (fan, pipe, hub, furnace…).

Questions, concerns and requests to Mission Support: –

Greenhab Report – February 19th

Here is the GreenHab Report


19 feb 2018

Environmental control:

(high winds today, door left closed, no heating)


Shade cloth on

Average temperatures:

Both thermometers are out of service

The temperature didn’t go above 25°C, the sky stayed pretty cloudy

Hours of supplemental light: from 8pm to 11pm

Changes to crops: none

Daily water usage for crops: around 15 gallons

Time(s) of watering for crops: 10:30am and 6pm

Narrative: I started to do an excel inventory of the plants in the GreenHab, there is a lot of different crops (with a majority of tomatoes and lettuce), a few plants might give some harvest before the end of our rotation.

Support/supplies needed: one thermometer with humidity measure

EVA Report – February 19th

Hello Graeme,

Here is our first EVA’s reports. We experienced a better weather than expected so that we were able to stay longer on the outside and visit the surroundings!

EVA Report:
EVA #1 Crew members: Louis Mangin (EVA leader), Jérémy Auclair, Alexandre Martin,

Location: Around the Hab 12S-518100E-4250700N
Near Pooh’s Corner: 12S-518900E-4250900N
Under Eastern Kissing Camel Ridge: 12S-518500E-4249700N
Top of Hab Ridge: 12S-517800E-4249700N
Zubrin’s head: 12S-519500E-4248100N

Time: departure at 2:00 p.m. Duration: 1 hour 50 minutes

– Find a location to deploy our two atmosphere study devices – First EVA logger in situ test
– Exploration and EVA training for newcomers
– First photos and videos shots

1:58 p.m.: EVA#1 left Airlock
2:01 p.m.: EVA left the Hab using Deimos & Opportunity
2:09 p.m.: Arrival on first place of interest: Pooh’s Corner 2:40 p.m.: Leaving Pooh’s Corner
2:50 p.m.: Arrival on second place of interest
3:00 p.m.: Climbing to the top of Hab Ridge (third place of interest) 3:25 p.m.: Getting back on the rovers
3:35 p.m.: Going to Zubrin’s head to shoot photos
3:50 p.m.: Getting back in the Hab


Louis Mangin
Crew 189 Commander

Commander Report – February 19th

Dear Graeme and mission support,

here the Crew 189 Commander Report 19Feb2018

Sol 1

Dear all,

I can say that I am already very proud of my team.

After the birthday celebration of my colleague Louis, yesterday evening, the whole crew spent a first good but quite cold night on Mars. This morning, we did our first physical training orchestrated by my amazing fellow crewmate Louis Mangin. It was a perfect preparation for our first morning in simulation. We were all busy this morning, everyone at his post: the LOAC experiment being fixed by Jérémy and the MegaARES experiment being assembled by Gabriel on the lower deck, the mission planning being determined upstairs after our first brainstorming ever on the Red Planet. With Jérémy, we spent some time, adjusting the spacesuit to put a camel bag with sugared water just in case of a hypoglycemia during his first EVA.

After only two days, the rehydrated food seems to be accepted and liked now (perhaps the birthday cake I cooked yesterday helped)! I was glad to discover that my fellow crewmember Benoit is a really good cooker! Let’s see what we will cook during our mission!

As I and Louis already took part in the simulation last year, Louis was our first EVA leader and I was Habcom in order to teach our new Martian crewmembers what we already know. With Louis, we gave a little briefing about the spacesuits and the EVA protocols but I am sure we will do better after some EVAs. The first EVA team went outside the Hab this afternoon and guess what? We managed to respect our time scheduling from the first EVA! The EVA team, led by Louis, managed to find a good spot for the MegaARES and LOAC experiments. We also tested the EVA Logger experiment of Louis and I am sure it will be a very helpful instrument during each EVA !

The EVA come back went well and they seemed already used to the equipment even if they experienced foggy helmets. After the EVA, we debriefed it to know the feelings of our new colleagues; each one worked on experiments and reports. The dinner will focus on the EVA debriefing and EVA objectives of tomorrow, with a new dry food gastronomic dish (I think the very nice smell of fresh bread I cooked will help again to make my crew like the dry food (tricky

As you can imagine, Sol 1 was especially encouraging! I cannot wait for tomorrow to explore again the red cliffs of Mars with this new amazing crew! After this first entire day in simulation, it is obvious that the whole crew is well installed on Mars! We will deploy our first experimentstomorrow, stay tuned!

Ad Astra!

Victoria Da-Poian,

Commander of the sportive Crew 189

Sol Summary – February 19th

Crew 189 Sol Summary Report 19 Feb 2018

Sol 1

Summary Title: The day of Firsts

Author: Victoria Da-Poian

Mission Status: All nominal

Sol Activity Summary: First physical training, first EVA. We started our day by the very famous physical training of the ISAE Supaero crews. After a nice morning working on our mission planning, fixing our experiments, and getting ready for our first EVA, we had a nice lunch with our first dried vegetables. After the lunch, the all crew worked on the EVA preparation to help the ones going outside and to learn the EVAs functioning. The EVA went fine and we will deploy our experiments tomorrow. Our first bread is baking.

Look Ahead Plan: Because of the wind planned for today, only one part of the crew went out for an EVA in the Hab surroundings to select a location to deploy our air sensors, MegaARES and LOAC. Tomorrow the other part of the crew will go to their first EVA to deploy our experiments.

Anomalies in work: Water leak in the Hab Lower deck, broken fans in spacesuits, water boiler broken

Weather: Windy, cold in the station

Crew Physical Status: Fine, already a bit tired from the first physical training.

EVA: First one

Reports to be file:





Greenhab Officer

Astronomer Officer


Support Requested: new water boiler (please)

Journalist Report – February 18th

Journalist report, 02/17/18 and 02/18/18: last moments on Earth

Louis had been the last to join us at Grand Junction, picked up by Gabriel, glad of his time of arrival. The mission was finally on its way! After a final brunch at the hotel, Victoria made a stock of bananas (which would be over even before we could see the Hab), it was time to leave civilization! The target: Hanksville, in the deepest paths of the Utah desert.

A journey not so comfortable at last, especially for Laurent, who carried the large MegaARES instrument that Gabriel left him. Nonetheless the road met its end, and we were getting deeper in the desert, towards the Hab. A few turns between giant red rocks later, we finally were able to see the station, among its little brothers, the laboratory, the GreenHab and the observatory. To welcome us, no previous crew was present contrary to the usual, but new stairs were being done, which will reduce falling risks and avoid injuries. It was then time for Shannon’s first briefing, completed with some advices from Attila the Peruvian, yet hard to understand because of Jack the dog, which seems to be into his teenager crisis.

Following a small cleaning session imposed by Victoria, it was now time for all of us to choose our rooms. Well, except for Gabriel, who inherited of the small bed under the celling, in order to take care of the hydraulic pump. A small rover/ATV trip later and we were taking our first meal in the station! A meal which finally revealed to be quite tasty: lyophilised ingredients are not so terrible. However, the crew started to feel really tired and everyone went to sleep to prepare the last pre-mission day!

A sleeping night which was short for Benoît and Louis, who had to bring back the location car to Grand Junction: it is an early morning for them! A few hours later, the remaining members of the crew tasted their first Martian breakfast with powdered orange juice and pancakes … powdered pancakes. The morning is used to put in place experiments which will be completed during the whole 3 weeks, and to finish some secondary tasks, like patch sewing, realized by Victoria, myself and Benoît for diverse results …

As both drivers came back from their journey and everyone filled their stomach, it was the time to take crew pictures! Despite some difficulties (one dog which repeatedly entered the field of view, and a tragic fall of the camera on the ground), photos were nonetheless satisfying. A last rover/ATV trip was organized, which proved to be quite hard due to continuous strong blowing winds and small dust storms that came with it, leading some members of the crew (especially Gabriel) to steal the look of a grandmother with his bandana.

The fatidic hour was near: while Shannon gave us her last advices, we prepared ourselves to the closing of the Hab door. Here begins 3 weeks without seeing the Sun without the helmet. As Shannon wished us good luck and left us, Gabriel finally grabbed the door handle to close the main door. After a few videos and pictures were taken by Jeremy, it was at last time to start the simulation.

6.05pm: the Hab door is closed, here begins the mission

Alexandre Martin, Crew Journalist MDRS 189.

Commander Report – February 18th

Crew 189 Commander Report 18Feb2018

Sol 0
Today began early for Benoit and I: because no crew was in the MDRS
before us to bring it back, we had to bring the rental car back to
Grand Junction. During this time, the rest of us had a bit of free
time to rest, then was trained by Dr Rupert on the Hab habits and map
use. We then had lunch together, and started the afternoon by our crew
official photo, in front of the hab. We then tried go on the ATVs but
quickly ended up in a little sand storm, so that we didn’t go very
After a last briefing more centered on simulation philosophy, came the
time to officially close the airlock. Even if it was less magical than
last year, it was still very symbolic: we are now by ourselves, and
for 3 weeks.
Everything is now slowly getting in order: everybody finds his role,
we wrote our first reports today, tested various equipment and look
forward to our first EVA, if the weather permits it! We want both to
train our new crewmembers about EVA specifications, and to find an
acceptable spot to display our experiments. As these will need to
remain on the field as long as possible, we want to make it a priority
for the beginning of the mission.
Everything is going pretty fast at the moment, but I think tomorrow
will be our last day of rush, and that after it, everybody will be
properly settled, and ready for the real business. Anyway, I see
excitement in all of us, and it is a good thing: everybody seems happy
to be here and motivated. I look forward to seeing what this crew is
going to accomplish!

Louis Mangin

Operations Report – February 18th

Crew 189 Operations Report 18Feb2018

SOL: 0

Name of person filing report: Gabriel PAYEN

Non-nominal systems: Pipe leak in low deck, water heater, two thermometers in the Greenhab.

Notes on non-nominal systems: –

Generator (hours run): 14 hours during the night (7 PM to 9 AM)

Solar— SOC % (Before generator is run at night) : 91%

Diesel – ?

Propane – 46.2 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons (estimation based the total height and capacity of the container)

Water (trailer) – 0 gallons.

Water (static) – 400 gallons (estimation based the total height and capacity of the tank)

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter (loft) : 0.44, measurement at 7 PM

Toilet tank emptied: No

ATV’s Used: All ATVs used except the one which doesn’t work (I forgot to check its number)

Oil Added : No

ATV Fuel Used: Gals

# Hours the ATVs were Used today: 20min

Notes on ATVs: –

Deimos rover used: Yes

Hours: 108.7

Beginning charge: ? (not checked)

Ending charge: 98%

Currently charging: No

Spirit rover used: No

Hours: 20.2

Beginning charge: –

Ending charge: –

Currently charging: Yes (Atilia may have unplugged Curiosity to plug in Spirit)

Opportunity rover used: Yes

Hours: 15.3

Beginning charge: ? (not checked)

Ending charge: 93% (100% after being charged)

Currently charging: No

Curiosity rover used: No

Hours: 12.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

HabCar used and why, where : Not used

General notes and comments: Last checks today with Shannon, some things to fix (see Non-nominal systems). Jérémy tried to run his aerosol counter device (LOAC) in the Hab but it doesn’t work. It turns on but doesn’t begin counting. The crew will try to fix it tomorrow.

Summary of internet: –

Summary of suits and radios: All air pipes had been tested and only one doesn’t work (backpack 3 – left one). All radios had been tested and work

Summary of Hab operations: –

Summary of GreenHab operations: –

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support:


The GreenHab officer would need a new thermometer.

-Pipe leaks

A bucket is beneath the pipe and receives a liquid which seems to be bad smelling water. Tape would stop the flow, but it only seems to be a temporary solution. Tomorrow I’ll try to find where the leak comes from.

-Water heater

Shannon seems to have fixed it before the very beginning of the SIM. We’ll see it when a crew member will take the first shower of the mission.


Some checks have been forgotten (beginning charge of the rovers, numbers of ATVs used…). Sorry for that, I’ll try not to do theses mistakes anymore.